# Empirical sandwich variance estimator for iterated conditional expectation g-computation

### v2

@@ 4 Simulation study
For estimation of the parameters of interest, unstratified and stratified ICE g-computation M-estimators in (3) and (4) under correct model specification were assessed. M-estimators were evaluated at five difference sample sizes, $n\in\{250,500,1000,2000,5000\}$, with 5000 iterations each. The following metrics were evaluated: bias, empirical standard error (ESE), average standard error (ASE), standard error ratio (SER), and 95% CI coverage [34]. Bias was defined as the mean of the estimated mean minus the true mean under the plan, with the true mean determined by simulating 10 million observations under the plan. ESE was estimated by the standard deviation of the point estimates of the simulation. ASE was estimated by the mean of the estimated standard errors. SER was defined as the ASE divided by the ESE. CI coverage was estimated by the proportion of 95% CI that contained the true mean under the plan. Whether the M-estimator procedure failed to find the root within 10000 iterations was also tracked. Failures to converge were ignored for computation of the other metrics.
@@ 4.1 Results
Simulation results are presented in Tables 2-3. As seen across sample sizes, both unstratified and stratified ICE g-computation M-estimators were approximately unbiased under correct model specification. In general, unstratified ICE g-computation has a smaller ESE, and thus was more precise, relative to stratified ICE g-computation. However, the difference between the ESE of the stratified and unstratified approaches diminished as sample sizes increased. For small sample sizes, like $n=250$, stratified ICE g-computation occasionally failed to converge (up to 6%). Both the differences in precision and failure to converge are likely due to data becoming sparse (i.e., few observations followed the plan of interest and thus the nuisance models either could not be estimated or had large variances). These issues were more pronounced for the always-treat parameter as treatment was less prevalent at all time points in the data generating mechanism.
@@ 5 Example
The parameters of interest for our analysis was the prevalence of hypertension at wave IV had none of the defined population been current cigarette smokers at waves I and III, ${\color[rgb]{0,0,0}\mu_{2}(0,0)}=\mathbb{E}[Y_{i,2}(0,0)]$. Additionally, we assessed the causal effect of this smoking ban relative to the natural course (i.e., observed patterns of smoking) [39], ${\color[rgb]{0,0,0}\delta}_{2}={\color[rgb]{0,0,0}\mu_{2}(0,0)}-{\color[rgb]{0,0,0}\mu_{2}}$ where ${\color[rgb]{0,0,0}\mu_{2}}=\mathbb{E}[Y_{i,2}]$. At wave IV, participants with type I (systolic between 140-159 or diastolic between 90-99) or type II (systolic 160+ or diastolic 100+) hypertension as measured by the interviewer were classified as having hypertension. For each visit, current cigarette smoking was defined as those who reported smoking cigarettes at least one day in the previous 30 days from the interview data. Those who reported zero days or never smoking were classified as not current cigarette smokers.
@@
For estimation of the mean under a smoking ban, the unstratified ICE g-computation estimator, ${\color[rgb]{0,0,0}\hat{\mu}_{2}(0,0)}$, was used. Rather than calculating the mean under the natural course directly, we instead used unstratified ICE g-computation with the observed values of $\bar{A}_{i,2}$. This approach has the advantage of accounting for informative loss to follow-up by $(\bar{A}_{i,k},\bar{L}_{i,k})$, unlike simply taking the mean of the observed $Y_{2}$. The ICE g-computation estimators for each plan were stacked together. Finally, an estimating function for the difference between the proportion under the smoking ban and natural course was stacked as well. The full set of stacking estimating equations for $\hat{\delta}_{2}$ were
$\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,2}=0)\left\{Y_{i,2}-\text{expit}(X_{i,1}^{T}\hat{\beta}_{1})\right\}X_{i,1}\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-{\color[rgb]{0,0,0}\hat{\mu}_{2}(0,0)}\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*^{\prime}}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*^{\prime}}-{\color[rgb]{0,0,0}\hat{\mu}_{2}}\\ ({\color[rgb]{0,0,0}\hat{\mu}_{2}(0,0)}-{\color[rgb]{0,0,0}\hat{\mu}_{2}})-\hat{\delta}_{2}\end{bmatrix}=\boldsymbol{0}$
@@
Analyses were conducted with Python 3.9.4 and replicated in R 4.2.0 (Vienna, Austria) with the geex, numDeriv, and rootSolve packages [20, 43, 44]. Data are freely available from the University of North Carolina at Chapel Hill Dataverse hosted by the Odum Institute [45, 46, 47]. Code to preprocess the data set and replicate the analysis is provided at gihtub.com/pzivich/publications-code. To compare runtimes, several implementations of ICE g-computation were considered. As a reference, ICE g-computation was implemented using successive generalized linear models, with the variance estimated using a nonparametric bootstrap with 500 resamples. Point estimates for nuisance models were computed using iteratively reweighted least squares with statsmodels [48]. Bootstrap iterations were run both in sequence and up to seven in parallel. Two implementations of M-estimators were compared. First, the solution $\hat{\boldsymbol{\theta}}$ was found via root-finding with the Levenberg-Marquardt with generic starting values (i.e., $0.18$ for the causal means and $0.0$ for all other parameters). As this approach involves the simultaneous estimation of many parameters, the root-finding procedure can take may iterations to converge. So, an implementation that first solves for the nuisance parameters using iteratively reweighted least squares and then uses Levenberg-Marquardt to solve for the parameters of interest was also considered. While the reported runtime still includes solving for the nuisance parameters subsets, the overall runtime is expected to be lower. Runtime results were reported for Python only.
@@ 5.1 Results
After application of the exclusion criteria, 5657 (87%) observations remained in the analytic data set. Descriptive statistics for the analytic data set are provided in Tables 4-5. Between waves I and III, 1694 (30%) observations were censored. Between waves III and IV, an additional 594 (15%) of observations were censored. Had all current smoking been prevented at waves I and III, the estimated prevalence of hypertension at wave IV would have been 0.175 (95% CI: 0.157, 0.193), which is 1.15 percentage points lower (95% CI: -0.025, 0.002) than the natural course. The sandwich variance estimator provided a similar CI to the bootstrap but was substantially faster (Table 6). This result remained true even when up to seven bootstrap iterations were run in parallel. The reduction in runtime was the greatest when the nuisance parameters were solved outside of the M-estimator, indicating most of the runtime for the M-estimator with generic starting values was spent finding the roots of the estimating equations. Altogether, this suggests that M-estimators with a bit more involved coding can greatly reduce runtimes, but even out-of-the-box implementations can still reduce runtimes.
@@ 6 Conclusions
Here, we framed the ICE g-computation estimator as an M-estimator to reduce the computational burden of variance estimation with the nonparametric bootstrap. Performance of the empirical sandwich variance estimator in the simulation study aligned with expectations and provided notable reductions in runtimes in the applied example. As indicated in our simulations, stratified ICE g-computation may fail to converge or have poor performance with small sample sizes. In these cases, unstratified ICE g-computation may be preferred, under the additional assumption that the parametric constraints used in the models are deemed to be close approximations.
This paper focused on deterministic plans that did not depend on the natural course. However, ICE g-computation has been extended for plans that depend on the natural course or stochastic plans (i.e., plans where treatment is assigned probabilistically) [50]. Generalizations of the proposed M-estimators could also be developed for these extensions of ICE g-computation. Use of M-estimators is also not limited to g-computation. Inverse probability weighting estimators for longitudinal data can be expressed as stacked estimating equations [9], thereby avoiding the conservative estimation of the variance via the ‘robust’ variance estimator in some settings [1, 51]. Multiply-robust estimators can also be expressed as estimating equations [6, 7, 9, 52, 53]. For some versions of these multiply-robust estimators, closed-form variance estimators based on the influence curve are available [52, 53]. For those without a simple or readily available closed-form expression for the variance, the empirical sandwich variance estimator remains an appealing option.
@@ Acknowledgments
$\sum_{i=1}^{n}\psi_{U}(O_{i};\hat{\boldsymbol{\theta}})=\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,\tau}=0,Y_{i,\tau-1}=0)\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\hat{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0,Y_{i,\tau-2}=0)\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\hat{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-{\color[rgb]{0,0,0}\mu_{\tau}(\bar{a}_{\tau-1}^{*})}\end{bmatrix}=\boldsymbol{0}$

### v3

@@ 4 Simulation study
For estimation of the parameters of interest, unstratified and stratified ICE g-computation estimating equations under correct model specification were assessed. Estimators were evaluated at five difference sample sizes, $n\in\{250,500,1000,2000,5000\}$, with 5000 iterations each. The following metrics were evaluated: bias, empirical standard error (ESE), average standard error (ASE), standard error ratio (SER), and 95% CI coverage [34]. Bias was defined as the mean of the estimated mean minus the true mean under the plan, with the true mean determined by simulating 10 million observations under the plan. ESE was estimated by the standard deviation of the point estimates of the simulation. ASE was estimated by the mean of the estimated standard errors. SER was defined as the ASE divided by the ESE. CI coverage was estimated by the proportion of 95% CI that contained the true mean under the plan. Whether the root-finding procedure failed within 10000 iterations was also tracked. Failures to converge were ignored for computation of the other metrics.
@@ 4.1 Results
Simulation results are presented in Tables 2-3. As seen across sample sizes, both unstratified and stratified ICE g-computation estimators were approximately unbiased under correct model specification. In general, unstratified ICE g-computation has a smaller ESE, and thus was more precise, relative to stratified ICE g-computation. However, the difference between the ESE of the stratified and unstratified approaches diminished as sample sizes increased. For small sample sizes, like $n=250$, stratified ICE g-computation occasionally failed to converge (up to 6%). Both the differences in precision and failure to converge are likely due to data becoming sparse (i.e., few observations followed the plan of interest and thus the nuisance models either could not be estimated or had large variances). These issues were more pronounced for the always-treat parameter as treatment was less prevalent at all time points in the data generating mechanism.
@@ 5 Example
The parameters of interest for our analysis was the prevalence of hypertension at wave IV had none of the defined population been current cigarette smokers at waves I and III, $\mu_{2}(0,0)=\mathbb{E}[Y_{i,2}(0,0)]$. Additionally, we assessed the causal effect of this smoking ban relative to the natural course (i.e., observed patterns of smoking) [39], $\delta_{2}=\mu_{2}(0,0)-\mu_{2}$ where $\mu_{2}=\mathbb{E}[Y_{i,2}]$. At wave IV, participants with type I (systolic between 140-159 or diastolic between 90-99) or type II (systolic 160+ or diastolic 100+) hypertension as measured by the interviewer were classified as having hypertension. For each visit, current cigarette smoking was defined as those who reported smoking cigarettes at least one day in the previous 30 days from the interview data. Those who reported zero days or never smoking were classified as not current cigarette smokers.
@@
For estimation of the mean under a smoking ban, the unstratified ICE g-computation estimator, $\hat{\mu}_{2}(0,0)$, was used. Rather than calculating the mean under the natural course directly, we instead used unstratified ICE g-computation with the observed values of $\bar{A}_{i,2}$. This approach has the advantage of accounting for informative loss to follow-up by $(\bar{A}_{i,k},\bar{L}_{i,k})$, unlike simply taking the mean of the observed $Y_{2}$. The ICE g-computation estimators for each plan were stacked together. Finally, an estimating function for the difference between the proportion under the smoking ban and natural course was stacked as well. The full set of stacking estimating equations for $\hat{\delta}_{2}$ were
$\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,2}=0)\left\{Y_{i,2}-\text{expit}(X_{i,1}^{T}\hat{\beta}_{1})\right\}X_{i,1}\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-\hat{\mu}_{2}(0,0)\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*^{\prime}}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*^{\prime}}-\hat{\mu}_{2}\\ (\hat{\mu}_{2}(0,0)-\hat{\mu}_{2})-\hat{\delta}_{2}\end{bmatrix}=\boldsymbol{0}$
@@
Analyses were conducted with Python 3.9.4 and replicated in R 4.2.0 (Vienna, Austria) with the geex, numDeriv, and rootSolve packages [19, 43, 44]. Data are freely available from the University of North Carolina at Chapel Hill Dataverse hosted by the Odum Institute [45, 46, 47]. Code to preprocess the data set and replicate the analysis is provided at gihtub.com/pzivich/publications-code. To compare runtimes, several implementations of ICE g-computation were considered. As a reference, ICE g-computation was implemented using successive generalized linear models, with the variance estimated using a nonparametric bootstrap with 500 resamples. Point estimates for nuisance models were computed using iteratively reweighted least squares with statsmodels [48]. Bootstrap iterations were run both in sequence and up to seven in parallel. Two implementations of the estimating equations were compared. First, the solution $\hat{\boldsymbol{\theta}}$ was found via root-finding with the Levenberg-Marquardt with generic starting values (i.e., $0.18$ for the causal means and $0.0$ for all other parameters). As this approach involves the simultaneous estimation of many parameters, the root-finding procedure can take may iterations to converge. So, an implementation that first solves for the nuisance parameters using iteratively reweighted least squares and then uses Levenberg-Marquardt to solve for the parameters of interest was also considered. While the reported runtime still includes solving for the nuisance parameters subsets, the overall runtime is expected to be lower. Runtime results were reported for Python only.
@@ 5.1 Results
After application of the exclusion criteria, 5657 (87%) observations remained in the analytic data set. Descriptive statistics for the analytic data set are provided in Tables 4-5. Between waves I and III, 1694 (30%) observations were censored. Between waves III and IV, an additional 594 (15%) of observations were censored. Had all current smoking been prevented at waves I and III, the estimated prevalence of hypertension at wave IV would have been 0.175 (95% CI: 0.157, 0.193), which is 1.15 percentage points lower (95% CI: -0.025, 0.002) than the natural course. The sandwich variance estimator provided a similar CI to the bootstrap but was substantially faster (Table 6). This result remained true even when up to seven bootstrap iterations were run in parallel. The reduction in runtime was the greatest when the nuisance parameters were solved prior to root-finding, indicating most of the runtime with generic starting values was spent finding the roots of the estimating equations. Altogether, this suggests that implementations with a bit more involved coding can greatly reduce runtimes, but even out-of-the-box implementations can still reduce runtimes.
@@ 6 Conclusions
Here, we expressed the ICE g-computation estimator as a set of estimating equations to reduce the computational burden of variance estimation with the nonparametric bootstrap. Performance of the empirical sandwich variance estimator in the simulation study aligned with expectations and provided notable reductions in runtimes in the applied example. As indicated by our simulations, stratified ICE g-computation may fail to converge or have poor performance with small sample sizes. In these cases, unstratified ICE g-computation may be preferred, under the additional assumption that the parametric constraints used in the models are deemed to be close approximations.
This paper focused on deterministic plans that did not depend on the natural course. However, ICE g-computation has been extended for plans that depend on the natural course or stochastic plans (i.e., plans where treatment is assigned probabilistically) [50]. Generalizations of the proposed estimating equations could also be developed for these extensions of ICE g-computation. Use of estimating equations is also not limited to g-computation. Inverse probability weighting estimators for longitudinal data can be expressed as stacked estimating equations [9], thereby avoiding the conservative estimation of the variance via the ‘robust’ variance estimator in some settings [1, 51]. Multiply-robust estimators can also be expressed as estimating equations [6, 7, 9, 52, 53]. For some versions of these multiply-robust estimators, closed-form variance estimators based on the influence curve are available [52, 53]. However, the empirical sandwich variance estimator may offer multiply-robust inference, unlike the influence curve variance estimators [54]. Therefore, the empirical sandwich variance estimator remains an appealing option.
@@ Acknowledgments
$\sum_{i=1}^{n}\psi_{U}(O_{i};\hat{\boldsymbol{\theta}})=\sum_{i=1}^{n}\begin{bmatrix}I(C_{i,\tau}=0,Y_{i,\tau-1}=0)\left\{Y_{i,\tau}-\text{expit}(X_{i,\tau-1}^{T}\hat{\beta}_{\tau-1})\right\}X_{i,\tau-1}\\ I(C_{i,\tau-1}=0,Y_{i,\tau-2}=0)\left\{\tilde{Y}_{i,\tau-1}^{*}-\text{expit}(X_{i,\tau-2}^{T}\hat{\beta}_{\tau-2})\right\}X_{i,\tau-2}\\ \vdots\\ I(C_{i,1}=0)\left\{\tilde{Y}_{i,1}^{*}-\text{expit}(X_{i,0}^{T}\hat{\beta}_{0})\right\}X_{i,0}\\ \tilde{Y}_{i,0}^{*}-\mu_{\tau}(\bar{a}_{\tau-1}^{*})\end{bmatrix}=\boldsymbol{0}$